OpenStudy (anonymous):

I AM AWARDING MEDALS! Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.)

OpenStudy (anonymous):

as the kids say "tmi"

OpenStudy (anonymous):

Hey its @julian25 , you gonna help

OpenStudy (anonymous):

Oh, sorry @sa

OpenStudy (anonymous):

@satellite73 *

OpenStudy (anonymous):

yes give a minute

OpenStudy (anonymous):

you can't?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

how do you think i got this nice green bike?

OpenStudy (anonymous):

auction

OpenStudy (anonymous):

lololol hahaha!

OpenStudy (anonymous):

there are many different ways you can get a solution here one way is to start by writing \[y=a(x+6)(x-2)\]

OpenStudy (anonymous):

Why do you do that

OpenStudy (anonymous):

then since you have \((-2,-16)\) on the graph you can solve for \(a\) by writing \[-16=(-2+6)(-2-2)\] and solve

OpenStudy (anonymous):

if you know the zeros, you know how it factors

OpenStudy (anonymous):

the other way it to write it in "vertex" form as \[y=a(x+2)^2-16\]

OpenStudy (anonymous):

y=a(x-h)^2+k (h,k) is the vertex and for x and y we can use one of the x intercepts. Let's choose (2,0) 0=a(2+2)^2-16 0=16a-16 a=1 y=1(x+2)^2-16 is the required equation. If you want it in a different form, it can be rearranged y=x^2+4x+4-16 y=x^2+4x-12

OpenStudy (anonymous):

again you can solve for \(a\) by putting \[0=a(2+2)^2-16\]

OpenStudy (anonymous):

ahh that is what @FriedRice did above

OpenStudy (anonymous):

as you can see you have too much information one method required only the vertex and one point the first method i wrote used both zeros, but ignored the fact that \((-2,-16)\) was the vertex

OpenStudy (anonymous):

But wait I don't g et what to do? @satellite73

OpenStudy (anonymous):

you are looking for a quadratic with three features 1) the vertex is \((-2,-16)\) 2) one of the zeros is \(x=2\) i.e. \((2,0)\) is on the graph 3) another zero is \(x=-6\) so \((-6,0)\) is on the graph

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can do it as @FriedRice did above write the quadratic in vertex form as \[y=a(x+2)^2-16\] and the only number you do not know is \(a\)

OpenStudy (anonymous):

but you know that if \(x=2\) you must have \(y=0\) because you are told that one of the zeros is \(2\)

OpenStudy (anonymous):

So 0=a(2+2)^2-16

OpenStudy (anonymous):

so replace \(y\) by \(0\) and \(x\) by \(2\) and get the equation \[0=a(2+2)^2-16\] solve for \(a\) via \[0=16a-16\\ 16=16a\\ 1=a\]

OpenStudy (anonymous):

yes, what you said

OpenStudy (anonymous):

So 0=0 if you work through

OpenStudy (anonymous):

now we know that \(a=1\) so it is \[y=(x+2)^2-16\] which you can expand if you like

OpenStudy (anonymous):

no, not \(0=0\) but rather \(a=1\)

OpenStudy (anonymous):

yes, watch my work 0=(2+2)^2-16 parentheses 0=4^2-16 exponents 0=16-16 subtraction 0=0

OpenStudy (anonymous):

@satellite73